The intensity distribution of solar radiation peaks at a wavelength of approximately 500 nm. Assuming the Sun radiates like a black body, calculate the flux of photons leaving the sun in W/m2.
1
Expert's answer
2020-07-27T07:23:33-0400

As per wein's displacement law

"\lambda_m(\mu m)=\frac{2900}{T}"

"T=\frac{2900}{500\times 10 ^{-3} \mu m}=5800 K"


Now, we know that

Solar irradiation ="H= \sigma T^4= 5.67\times 10^{-12}\times (5800^4)"

H= 6416.45 "\frac{W}{m^2}"